# Shock waves in nonlinear transmission lines

### v3

@@ Abstract
We consider interaction between the small amplitude travelling waves (”sound”) and the shock waves in the transmission line containing both nonlinear capacitors and nonlinear inductors. We calculate for the ”sound” wave the coefficient of reflection from (the coefficient of transmission through) the shock wave. These coefficients are expressed in terms of the wave speeds and the wave impedances. When only the capacitors or only the inductors are nonlinear, the coefficients are expressed in terms of the wave speeds only. We explicitly include into consideration of the shocks the dissipation, introducing ohmic resistors shunting the inductors and also in series with the capacitors. This allows us to describe the shocks as physical objects of finite width and study their profiles. In some particular cases the profiles were obtained in terms of elementary functions.
@@ I Introduction
In the present paper we consider the travelling waves in the transmission line where, in the general case, both the inductors and the capacitors are nonlinear. In the first part of the paper, generalizing our previous results kogan4 ; kogan5 , we calculate the ”sound” reflection from and transmission through the shock wave in these nonlinear transmission lines. In the second part of the paper we include explicitly the ohmic losses (which are of course necessary for the existence of the shock waves) into the circuit equations. This allows us to study the profile of the shock wave.
@@ II The circuit equations
We take the voltages on the capacitors $V_{n}$ and the current through the inductors $I_{n}$ as the dynamical variables. The circuit equations (Kirchhoff laws) are
@@
$\frac{d\Phi_{n+1}}{dt}$ $=V_{n}-V_{n+1},$ (1b)
@@
$Q_{n}$ is the charge of the nonlinear capacitor and ${\cal V}_{n}$ is the voltage on the nonlinear inductor.
In the continuum approximation we treat $n$ as the continuous variable $x$ and approximate the finite differences in the r.h.s. of the equations by the first derivatives with respect to $x$, after which the equations take the form
$\frac{\partial Q}{\partial t}$ $=-\frac{\partial I}{\partial x},$ (3a)
$\frac{\partial\Phi}{\partial t}$ $=-\frac{\partial V}{\partial x}$ (3b)
(we measure distance in the units of the transmission line period). To close the system (3) we should specify
$V$ $=V(Q),$ (4a)
$I$ $=\Phi(\Phi).$ (4b)
Further on we’ll consider $V$ and $I$ as known functions of $Q$ and $\Phi$ respectively.
III The ”sound” and the shocks
Because (3) is nonlinear, it can have a lot of different types of solutions. In this paper we’ll be interested in only two types of those. First type - small amplitude waves on a homogeneous background $\varphi_{0},V_{0}$.
$V$ $=V_{0}+v,$ (5a)
$I$ $=I_{0}+i.$ (5b)
For such waves Eq. (3) is simplified to
$C_{0}\frac{\partial v}{\partial t}$ $=-\frac{\partial i}{\partial x},$ (6a)
$L_{0}\frac{\partial i}{\partial t}$ $=-\frac{\partial v}{\partial x},$ (6b)
@@ III The ”sound” and the shocks
$C_{0}$ $=\left.\frac{dQ}{dV}\right|_{Q=Q_{0}},$ (7a)
$L_{0}$ $=\left.\frac{d\Phi}{dI}\right|_{\Phi=\Phi_{0}}.$ (7b)
The solutions of Eq. (6) (which, for brevity we’ll call the ”sound”) are right- and left-propagating travelling waves, each depending upon the single variable $\tau_{\pm}=t\mp x/u_{0}$, where
$u_{0}^{2}=\frac{1}{L_{0}C_{0}}$ (8)
is the ”sound” speed. The current and the voltage in the waves are connected by the equations
$v=\pm Z_{0}i,$ (9)
where
$Z_{0}=\sqrt{\frac{L_{0}}{C_{0}}}$ (10)
is the linear impedance of the transmission line.
@@
where
$V_{a,b}\equiv V(Q_{a,b}),\hskip 28.45274ptI_{a,b}\equiv I(\Phi_{a,b}),$ (12)
and $U_{21}$ is the shock wave speed. The indices 1 and 2, which refer to the quantities before and after the shock respectively, enter in the system (11) in a symmetrical way. However, in Section V we’ll show that the shock always propagates in such a way that whitham ; kogan4 ; kogan5
@@ IV The ”sound” reflection and transmission
$\left(\begin{array}[]{c|c}V_{1}&V_{2}\\ I_{1}&I_{2}\end{array}\right)=\left(\begin{array}[]{c|c}V_{b}&V_{a}+v^{(in)}+v^{(r)}\\ I_{b}&I_{a}+i^{(in)}+i^{(r)}\end{array}\right),$ (18)
@@
$U_{21}=U_{ab}+\delta U^{(refl)}.$ (19)
Substituting (18) into (11), in linear with respect to the ”sound” wave amplitudes approximation we obtain
$\delta U^{(refl)}\left(Q_{a}-Q_{b}\right)+U_{ab}C_{a}\left(v^{(in)}+v^{(r)}\right)=i^{(in)}+i^{(r)},$ (20a)
$\delta U^{(refl)}\left(\Phi_{a}-\Phi_{b}\right)+U_{ab}L_{a}\left(i^{(in)}+i^{(r)}\right)=v^{(in)}+v^{(r)}.$ (20b)
Dividing the equations (20) to each other we obtain
$u_{in}v^{(in)}-u_{r}v^{(r)}=Z_{ab}\left(u_{in}i^{(in)}-u_{r}i^{(r)}\right),$ (21)
@@
$Z_{ab}\equiv\frac{V_{a}-V_{b}}{I_{a}-I_{b}}=\frac{\Phi_{a}-\Phi_{b}}{Q_{a}-Q_{b}}.$ (22)
Taking into account the relations
$v^{(in)}$ $=Z_{a}i^{(in)},$ (23a)
$v^{(r)}$ $=-Z_{a}i^{(r)}$ (23b)
(the difference in the signs is because of the opposite directions of propagation of the two waves) and solving (21) we obtain for the reflection coefficient
$R\equiv\frac{i^{(r)}}{i^{(in)}}=\frac{\left(Z_{ab}-Z_{a}\right)u_{in}}{\left(Z_{ab}+Z_{a}\right)u_{r}}.$ (24)
@@
$\left(\begin{array}[]{c|c}V_{1}&V_{2}\\ I_{1}&I_{2}\end{array}\right)=\left(\begin{array}[]{c|c}V_{b}+V^{(in)}&V_{a}+v^{(t)}\\ I_{b}+i^{(in)}&I_{a}+i^{(t)}\end{array}\right),$ (29)
@@
$\delta U^{(trans)}\left(Q_{a}-Q_{b}\right)+U_{ab}\left(C_{a}v^{(t)}-C_{b}v^{(in)}\right)=i^{(t)}-i^{(in)},$ (30a)
$\delta U^{(trans)}\left(\Phi_{a}-\Phi_{b}\right)+U_{ab}\left(L_{a}i^{(t)}-L_{b}i^{(in)}\right)=v^{(t)}-v^{(in)}.$ (30b)
Dividing the equations (30) to each other we obtain
$u_{a}u_{in}v^{(in)}-u_{b}u_{t}v^{(t)}=Z_{ab}\left(u_{a}u_{in}i^{(in)}-u_{b}u_{t}i^{(t)}\right),$
where $u_{in}=u_{b}+U_{ab}$ is the speed of the incident ”sound” wave relative to the shock wave, and $u_{t}=u_{a}+U_{ab}$ is the speed of the transmitted ”sound” wave relative to the shock wave. Taking into account the relations
$v^{(in)}$ $=-Z_{b}i^{(in)},$ (32a)
$v^{(t)}$ $=-Z_{a}i^{(t)}$ (32b)
and solving (IV) we obtain for the transmission coefficient
$T\equiv\frac{i^{(t)}}{i^{(in)}}=\frac{\left(Z_{ab}+Z_{b}\right)u_{a}u_{in}}{\left(Z_{ab}+Z_{a}\right)u_{b}u_{t}}.$ (33)
As one could have expected, the modulus of the ”sound” reflection coefficient is less than one, and it goes to zero when the intensity of the shock wave decreases, that is when $Q_{a}\to Q_{b},\Phi_{a}\to\Phi_{b}$, in other words, when the shock wave itself is nearly the ”sound” wave. Similarly, in these circumstances the transmission coefficient goes to 1.
Looking back at (21) and (IV) we wonder whether these equations could have been written down by considering propagation of the waves in the frame of reference moving with the shock wave.
@@ IV.1.1 Linear capacitor
For linear capacitor ($Q=VC$, where $C$ is a constant) from (8) and (10) follows that
$Z_{a,b}=\frac{1}{Cu_{a,b}},$ (34)
and from (11) and (22) follows that
$Z_{ab}=\frac{1}{CU_{ab}}.$ (35)
@@
$R$ $=\frac{u_{in}^{2}}{u_{r}^{2}},$ (36a)
$T$ $=\frac{u_{a}^{2}u_{in}^{2}}{u_{b}^{2}u_{t}^{2}},$ (36b)
and from (8) and (11) follows
@@
Consider the case of JTL kogan4 . In this case from the second Josephson law
${\cal V}=\frac{\hbar}{2e}\frac{d\varphi}{dt},$ (38)
where $\varphi$ is Josephson phase, we obtain
$\Phi=\frac{\hbar}{2e}\varphi.$ (39)
Hence the first Josephson law gives
@@
and we can easily calculate all the velocities. Equations (36) for the case of JTL were obtained in our previous publication kogan4 . (Comparing the results of Ref. kogan4 with those presented in this paper, one should pay attention to a slightly different definitions of the transmission and the reflection coefficients in the two papers.)
@@ IV.1.2 Linear inductor
For linear inductor ($\Phi=LI$, where $L=$ is a constant) from (8) and (10) follows that
$Z_{a,b}=Lu_{a,b},$ (41)
and from (11) and (22) follows that
$Z_{ab}=LU_{ab}.$ (42)
@@
$R$ $=-\frac{u_{in}^{2}}{u_{r}^{2}},$ (43a)
$T$ $=\frac{u_{a}u_{in}^{2}}{u_{b}u_{t}^{2}},$ (43b)
and from (8) and (11) follows
@@
Note that because of the symmetry of Eq. (3) with respect to simultaneous interchange
$Q$ $\leftrightarrows\Phi,$ (45a)
$V$ $\leftrightarrows I,$ (45b)
Eqs. (43), (44) follow from Eqs. (36), (37).
@@ V Lossy transmission line
From the five parameters ($Q_{1},Q_{2},\Phi_{1},\Phi_{2},U$) - three are arbitrary constants, and the other two are defined by the equations.
Both $(Q_{1},\Phi_{1})$ and $(Q_{2},\Phi_{2})$ are the fixed points (in the $(Q,\Phi)$ phase space), hence the r.h.s. of (48b) is equal to zero in both points. Equating these two values we recover the jump conditions (11b). Doing the same procedure with (48a) we recover the jump conditions (11a).
$(Q_{1},\Phi_{1})$ is the semistable fixed point of the system (48), and $(Q_{2},\Phi_{2})$ is the stable fixed point. (Other options would have demanded fine tuning of the parameters.) This means
@@ V.1 Half-nonlinear transmission line
Further analysis of the fully nonlinear line we relegate to Appendix B) and here consider transmission line with the linear capacitor. Because of the symmetry of Eq. (48) given by (45) (to which we should add $1/R_{L}\leftrightarrows R_{C}$), the case of the transmission line with linear capacitors and nonlinear inductors (JTL first and foremost) would be identical to the one considered in this Section..
Because of the relation $\Phi=LI$, where $L$ is a constant, we can exclude the current from (48) and after a bit of algebra obtain
@@ Acknowledgements.
We are grateful to A. M. Kamchatnov for sending us the manuscript of his (to be published) excellent book ”Theory of Nonlinear Waves”, which motivated the continuation of our previous studies. We are very grateful to J. Cuevas-Maraver, M. Goldstein, A. Kamchatnov, E. Konig, B. Malomed, A. Novic-Cohen and V. Verbuk for the insightful discussions. We are also grateful to the organizers of Nor-Amberd School in Theoretical Physics, the scientific atmosphere of which stimulated the presented research. The paper was finalized during the Author’s participation in School and Workshop on Dynamical Systems, ICTP, Trieste. The Author is grateful to the Center for the hospitality.
@@ A.1 Integration in terms of elementary functions
Hence the condition on the parameters of the problem, for (78) to be the solution of (77) is given by (71), and the constant multiplier in (78) should is given by (73). Substituting (78) into Eq. (74) we obtain the solution (72).
@@
If ${\cal I}$ is a polynomial function of $Q$, it should be a 3/2 order polynomial. And the boundary conditions (76) tell us that this polynomial should zero out at $Q=Q_{1,2}$.
Let us split in two one multiplier in the solution (78)
@@
Substituting the ansatz into (80) we see that it is in fact the solution of the equation, provided the parameters of the problem satisfy the relation (66) and $\chi$ is given by (68). Substituting thus obtained ${\cal I}(Q)$ into (74) we obtain (67)

### v4

@@ Abstract
In the first half of the paper we consider interaction between the small amplitude travelling waves (”sound”) and the shock waves in the transmission line containing both nonlinear capacitors and nonlinear inductors. We calculate the ”sound” wave coefficient of reflection from (coefficient of transmission through) the shock wave. These coefficients are expressed in terms of the speeds of the ”sound” waves relative to the shock and the wave impedances. In the second half of the paper we explicitly include into consideration the dissipation in the system, introducing ohmic resistors shunting the inductors and also in series with the capacitors. This allows us to justify the conditions on the shocks, postulated in the first half of the paper. This also allows us to describe the shocks as physical objects of finite width and study their profiles. In some particular cases the profiles were obtained in terms of elementary functions.
@@ I Introduction
In the present paper we consider the travelling waves in the transmission line where, in the general case, both the inductors and the capacitors are nonlinear. In the first part of the paper, generalizing our previous results kogan4 ; kogan5 , we calculate the ”sound” reflection from and transmission through the shock wave in these nonlinear transmission lines. In the second part of the paper we include explicitly the ohmic losses (which are of course necessary for the existence of the shock waves) into the circuit equations.
@@ II The circuit equations
We take the capacitors charges $Q_{n}$ and the integrated voltages on the inductors
[…]
as the dynamical variables. The circuit equations (Kirchhoff laws) are
@@
$\frac{d\Phi_{n+1}}{dt}$ $=V_{n}-V_{n+1}.$ (2b)
To close the system (2) we should specify the connection between the voltages on the capacitors $V_{n}$ and the charges and also between the currents through the inductors $I_{n}$ and the integrated voltages
$V_{n}$ $=V(Q_{n}),$ (3a)
$I_{n}$ $=I(\Phi_{n}).$ (3b)
Further on we’ll consider $V(Q)$ and $I(\Phi)$ as known functions.
In the continuum approximation we treat $n$ as the continuous variable $x$ (we measure distance in the units of the transmission line period) and approximate the finite differences in the r.h.s. of the equations by the first derivatives with respect to $x$, after which the equations take the form
$\frac{\partial Q}{\partial t}$ $=-\frac{\partial I}{\partial x},$ (4a)
$\frac{\partial\Phi}{\partial t}$ $=-\frac{\partial V}{\partial x}.$ (4b)
III The ”sound” and the shocks
Because (4) is nonlinear, it can have a lot of different types of solutions. In this paper we’ll be interested in only two types of those. First type - small amplitude waves on a homogeneous background $Q_{0},\Phi_{0}$.
$Q$ $=Q_{0}+q,$ (5a)
$\Phi$ $=\Phi_{0}+\phi.$ (5b)
For such waves Eq. (4) is simplified to
$\frac{\partial q}{\partial t}$ $=-\frac{1}{L_{0}}\frac{\partial\phi}{\partial x},$ (6a)
$\frac{\partial\phi}{\partial t}$ $=-\frac{1}{C_{0}}\frac{\partial q}{\partial x},$ (6b)
@@ III The ”sound” and the shocks
The solutions of Eq. (6) (which, for brevity we’ll call the ”sound”) are right- and left-propagating travelling waves, each depending upon the single variable $\tau_{\pm}=t\mp x/u_{0}$. The variables in the ”sound” wave are connected by the equations
$v$ =±$=\pm u_{0}\phi,$ (8a)
$i$ =±$=\pm u_{0}q,$ (8b)
$\phi$ =±$=\pm Z_{0}q,$ (8c)
@@
$u_{0}$ $=\frac{1}{\sqrt{L_{0}C_{0}}},$ (9a)
$Z_{0}$ $=\sqrt{\frac{L_{0}}{C_{0}}}$ (9b)
are the ”sound” speed and the linear impedance of the transmission line respectively.
@@
where $V_{a,b}\equiv V(Q_{a,b})$, $I_{a,b}\equiv I(\Phi_{a,b})$, and $U_{21}$ is the shock wave speed. The indices 1 and 2, which refer to the quantities before and after the shock respectively, enter in the system (10) in a symmetrical way. However, in Section V we’ll show that the shock always propagates in such a way that whitham ; kogan4 ; kogan5
@@ IV The ”sound” reflection and transmission
$\left(\begin{array}[]{c|c}Q_{1}&Q_{2}\\ \Phi_{1}&\Phi_{2}\end{array}\right)=\left(\begin{array}[]{c|c}Q_{b}&Q_{a}+q^{(in)}+q^{(r)}\\ \Phi_{b}&\Phi_{a}+\phi^{(in)}+\phi^{(r)}\end{array}\right),$ (16)
@@
$U_{21}=U_{ab}+\delta U$ (17)
(same in the next problem). Substituting (16) into (10), in linear with respect to the ”sound” wave amplitudes approximation we obtain (taking into account (8a) and (8b))
$\delta U\left(Q_{a}-Q_{b}\right)+U_{ab}\left(q^{(r)}+q^{(in)}\right)=u_{a}\left(q^{(in)}-q^{(r)}\right),$ (18a)
$\delta U\left(\Phi_{a}-\Phi_{b}\right)+U_{ab}\left(\phi^{(r)}+\phi^{(in)}\right)=u_{a}\left(\phi^{(in)}-\phi^{(r)}\right).$ (18b)
Excluding $\delta U$ from (18) and taking into account (8c) we obtain for the reflection coefficient
$R\equiv\frac{\phi^{(r)}}{\phi^{(in)}}=\frac{\left(Z_{a}-Z_{ab}\right)u_{in}}{\left(Z_{a}+Z_{ab}\right)u_{r}},$ (19)
@@
$Z_{ab}\equiv\frac{\Phi_{a}-\Phi_{b}}{Q_{a}-Q_{b}}.$ (20)
Note that from (10) follows
$\frac{\Phi_{a}-\Phi_{b}}{Q_{a}-Q_{b}}=\frac{V_{a}-V_{b}}{I_{a}-I_{b}}.$ (21)
@@
$\left(\begin{array}[]{c|c}Q_{1}&Q_{2}\\ \Phi_{1}&\Phi_{2}\end{array}\right)=\left(\begin{array}[]{c|c}Q_{b}+q^{(in)}&Q_{a}+q^{(t)}\\ \Phi_{b}+\phi^{(in)}&\Phi_{a}+\phi^{(t)}\end{array}\right),$ (26)
@@
$\delta U\left(Q_{a}-Q_{b}\right)+U_{ab}\left(q^{(t)}-q^{(in)}\right)=u_{b}q^{(in)}-u_{a}q^{(t)},$ (27a)
$\delta U\left(\Phi_{a}-\Phi_{b}\right)+U_{ab}\left(\phi^{(t)}-\phi^{(in)}\right)=u_{b}\phi^{(in)}-u_{a}\phi^{(t)}.$ (27b)
Excluding $\delta U$ from (27) and taking into account (8c) we obtain for the transmission coefficient
$T\equiv\frac{\phi^{(t)}}{\phi^{(in)}}=\frac{\left(Z_{b}+Z_{ab}\right)Z_{a}u_{in}}{\left(Z_{a}+Z_{ab}\right)Z_{b}u_{t}},$ (28)
where $u_{in}=u_{b}+U_{ab}$ is the speed of the incident ”sound” wave relative to the shock wave, and $u_{t}=u_{a}+U_{ab}$ is the speed of the transmitted ”sound” wave relative to the shock wave.
As one could have expected, the modulus of the ”sound” reflection coefficient is less than one, and it goes to zero when the intensity of the shock wave decreases, that is when $Q_{a}\to Q_{b},\Phi_{a}\to\Phi_{b}$, in other words, when the shock wave itself is nearly the ”sound” wave. Similarly, in these circumstances the transmission coefficient goes to 1. Looking at (19) and (28) we wonder whether these equations could have been obtained by considering propagation of the waves in the frame of reference moving with the shock wave.
@@ IV.1.1 Linear capacitor
For linear capacitor ($V=Q/C$, where $C$ is a constant)
$Z_{a,b}=\frac{1}{Cu_{a,b}},\hskip 28.45274ptZ_{ab}=\frac{1}{CU_{ab}}.$ (29)
@@
$R$ $=-\frac{u_{in}^{2}}{u_{r}^{2}},$ (30a)
$T$ $=\frac{u_{in}^{2}}{u_{t}^{2}},$ (30b)
and from (9a) and (10) follows
@@
For example, for the JTL
@@
and all the velocities can be easily calculated. Equations (30) in this case were obtained in our previous publication kogan4 .
@@ IV.1.2 Linear inductor
For linear inductor ($I=\Phi/L$, where $L$ is a constant)
$Z_{a,b}=Lu_{a,b},\hskip 28.45274ptZ_{ab}=LU_{ab}.$ (33)
@@
$R$ $=\frac{u_{in}^{2}}{u_{r}^{2}},$ (34a)
$T$ $=\frac{Z_{a}u_{in}^{2}}{Z_{b}u_{t}^{2}},$ (34b)
and from (9a) and (10) follows
@@ V Lossy transmission line
Equating the r.h.s. of (38a) at $\tau=\pm\infty$, we recover the jump conditions (10a). Doing the same procedure with (38b) we recover the jump conditions (10b). Because the quantities ($Q_{1},Q_{2},\Phi_{1},\Phi_{2},U$), satisfy the constraint (21), Eq. (39) specifies 3 boundary conditions, which thus define also the shock speed $U$. Continuing the analysis of the boundary conditions, we notice that $(Q_{1},\Phi_{1})$ is the semistable fixed point of the system (38), and $(Q_{2},\Phi_{2})$ is the stable fixed point. (Other options would have demanded fine tuning of the parameters.) This means
@@ V.1 Half-nonlinear transmission line
Further analysis of the fully nonlinear line we relegate to Appendix B) and here consider transmission line with the linear capacitor. Because of the symmetry of Eq. (38) the case of the transmission line with linear capacitors and nonlinear inductors (JTL first and foremost) would be identical to the one considered in this Section.. We can exclude the current from (38) and after a bit of algebra obtain
@@ Acknowledgements.
We are grateful to A. M. Kamchatnov for sending us the manuscript of his (to be published) excellent book ”Theory of Nonlinear Waves”, which motivated the continuation of our previous studies. We are very grateful to J. Cuevas-Maraver, M. Goldstein, A. Kamchatnov, E. Konig, B. Malomed, A. Novic-Cohen and V. Verbuk for the insightful discussions. We are also grateful to the organizers of Nor-Amberd School in Theoretical Physics, the scientific atmosphere of which stimulated the presented research. The paper was finalized during the Author’s participation in School and Workshop on Dynamical Systems and Workshop on Localization and Ergodicity, ICTP, Trieste. The Author is grateful to the Center for the hospitality.
@@ A.1 Integration in terms of elementary functions
Hence the condition on the parameters of the problem, for (68) to be the solution of (67) is given by (61), and the constant multiplier in (68) should is given by (63). Substituting (68) into Eq. (64) and integrating we obtain the solution (62).
@@
If ${\cal I}$ is a polynomial function of $Q$, it should be a 3/2 order polynomial equal to zero at $Q=Q_{1,2}$. So let us split in two one multiplier in the solution (68)
@@
Substituting the ansatz into (70) we see that it is in fact the solution of the equation, provided the parameters of the problem satisfy the relation (56) and $\chi$ is given by (58). Substituting thus obtained ${\cal I}(Q)$ into (64) and integrating we obtain (57)